OpenStudy (baby456):

You run buisness making birdhouses.you spend 600 dollars to start and it cost 5 dollars to make each birdhouse.total cost to build 1,2,3,4,5,6,7 number of birdhoses.

OpenStudy (studygurl14):

To build one bird house: 600 to start + 5 for the house = 65$ 2 birdhouses: 600+ 5(2) = 600+5+5 = 600+10 = 610

OpenStudy (studygurl14):

Just keep going